My Weight is 74 kg.

2. Determine your weight (e.g., the force of gravity on you).
3. Determine what normal force would be required to accelerate you upwards from ground level at a rate of 2.0 or 3.0 m/s^2 (choose one acceleration value).
4. What would the normal force be if the acceleration described in (3) was directed downwards?
1
Expert's answer
2020-10-30T13:19:54-0400

Determine your weight (e.g., the force of gravity on you):


"W=F_G=mg=74\cdot9.8=725\text{ N}."


The normal force that would be required to accelerate you upwards from ground level at a rate of 2.0 m/s/s can be found from Newton's second law:


"N-mg=ma,\\
N=m(g+a)=74(9.8+2)=873\text{ N}."


If the acceleration described in the previous point was directed downwards, the normal force would be

"N-mg=-ma,\\
N=m(g-a)=74(9.8-2)=577\text{ N}."